# The human-machine paradox:

### v5

@@ 6 Discussion
Finally, to better understand the importance of each framework’s variable on bottom-line utility, we leverage introspection analysis on meta-models learned on our synthetic data by means of i) Permutation Importance analysis and ii) SHapley Additive exPlanations method (SHAP). Permutation Importance analysis (Breiman, 2001) is a computational method measuring the importance of a feature by determining the increase of the prediction error after the values of that feature are randomly permuted (a larger increase in error denotes a more important feature). In the context of our work, we apply Permutation Importance analysis on each of the models trained in cross validation, finally computing mean and standard deviation of the results; also, we aggregate categorical values encoded in one-hot variables in a unique feature, improving interpretability. In contrast, the SHAP method (Štrumbelj and Kononenko, 2014) (Datta et al., 2016) (Lundberg and Lee, 2017), named by the acronym of SHapley Additive exPlanations, computes the marginal contribution provided by each feature to the output for any given example, providing local interpretability. Then, global model explanations are obtained by aggregating Shapley values across all examples in the data set, suggesting which features are most influential overall, and how they impact the model’s predictions. As the features $c$ and $a$ have a strong correlation by construction, to avoid compromising the response of permutation analysis and SHAP (Strobl et al., 2008) (Lundberg and Lee, 2017), we combine these features in a unique new variable called c_ac\_a, which jointly represents skill policy and augmenting effects. According to these provisions, we learn two models: $\hat{f}:(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})\rightarrow\mu(U_{c,d})$ and $\hat{f}:(c\_a,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})\rightarrow\mu(U_{c})$. The absence of correlation between each models’ variable is verified by computing their adjusted Generalized Variance Inflation Factor (Adj.GVIF) (Fox and Monette, 1992), which results consistently in the range $[1.00,1.01]$ (no correlation). With regards to the modeling strategy, we implement Random Forest regressors (RFr) and Gradient Boosting regressors (GBr), assessing their performance through 10-Fold Cross Validation (metrics: $MSE$ and $R^{2}$, see Table 4; the purpose of measuring the predictive performance is only to make sure that models are effectively compressing data and learning the information encoded in it, an essential pre-requisite for the interpretation of variable importance). The results illustrated in Figure 14 and in Figure 15 suggest that the skill/augmentation policy c_ac\_a is the most important variable in terms of effect on bottom-line utility. Naturally, towards $\mu(U^{s}_{[c],[d]})$, also task difficulty $d$ is a remarkably important variable, increasingly when the distributions $P(\theta)_{c,d,e}$ differ upon changes in $d$ (the setup of the experiments is discriminant for the emergence of this effect). Then, a group of variables emerges for a second-level importance, including the performance threshold triggering the error shocks ($t_{err}$), the cost of errors ($c_{err}$), the baseline margin of contribution associated to the human-machine skill policy ($mc_{HM}$) and the performance-to-output function $g^{\prime}$. In contrast, the augmenting factor ($\gamma_{HM}$) shows to play a minor role, which suggests that the augmenting gain may be less important than the dynamics of skill collaboration and interplay.
@@
Table 4: Combined results of the 10-Fold Cross Validation (MSE, $R^{2}$) for Exp. 1 and Exp. 2
Exp. 1 Exp. 2
RFr GBr RFr GBr
Meta-model MSE $R^{2}$ MSE $R^{2}$ MSE $R^{2}$ MSE $R^{2}$
$\mu(U_{c,d})=\hat{f}(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$ 0.000 0.994 0.004 0.932 0.000 0.996 0.003 0.945
$\mu(U_{c})=\hat{f}(c\_a,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$ 0.000 0.991 0.002 0.925 0.000 0.996 0.002 0.957
(a) Exp. 1, RFr
(b) Exp. 1, GBr
(c) Exp. 2, RFr
(d) Exp. 2, GBr
(e) Exp. 1, RFr
(f) Exp. 1, GBr
(g) Exp. 2, RFr
(h) Exp. 2, GBr
Figure 14: Aggregated Permutation Importance analysis. The first row shows the mean decrease of $R^{2}$ for the model $\mu(U_{c,d})=\hat{f}(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$, while the second row relates to the model $\mu(U_{c})=\hat{f}(c\_a,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$.
(a) Exp. 1, RFr
(b) Exp. 1, GBr
(c) Exp. 2, RFr
(d) Exp. 2, GBr
(e) Exp. 1, RFr
(f) Exp. 1, GBr
(g) Exp. 2, RFr
(h) Exp. 2, GBr
@@
Figure 15: SHAP summary (beeswarm) plot illustrating variables’ impact on the output of models. The first row shows the mean decrease of $R^{2}$ for the model $\mu(U_{c,d})=\hat{f}(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$, while the second row relates to the model $\mu(U_{c})=\hat{f}(c\_a,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$. Variables are ranked by global importance. Each point represents the SHAP value for a specific instance, with the color indicating the feature’s value. Categorical values (extremes): for task difficulty $d$, blue dots correspond to Low and red dots to High; for skill/augmentation policy c_ac\_a, blue dots correspond to H individual and red dots to HM superpower, with intermediate color scale passing through HMHM mean, $M$ individual (grey), and HMHM collaborate; for the function $g^{\prime}$, the blue dots represent logistic, while red dots represent inverse_logistic.
@@ 7 Limitations
Combining Monte Carlo simulations with machine learning meta-models can provide valuable insights, but limitations must be acknowledged to contextualize findings appropriately. The primary limitation of this study is that its conclusions are derived from synthetically generated data. Essentially, the simulation creates an idealized in silico environment whose potential gaps with the real-world should always be considered with the highest regard, as complex or unmodeled interactions could fundamentally alter the outcomes (Law et al., 2007). As a consequence, the validity of our findings is contingent upon the fidelity of our model, which is governed by two sets of assumptions: design hypotheses, related to the theoretical mechanics and interactions encoded in the model, and setup hypotheses, concerning the specific characterization of the scenarios of interest through parameters and probability distributions chosen for the stochastic elements. Then, before verification and validation, results and insights should be intended as specific to the explored parameter space. Design hypotheses include the specific choice in terms of distributions (for example, the choice of $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$) and the logics of computation of economic measures (for example, the cost of errors). Another arbitrary design choice pertains the specific integration of stochastic effects, which in our work are limited to the task-execution performance, leaving the other measures of economic impact to deterministic computation. This can potentially inject a re-scaling effect down through the funnel of economic value, thus making the dynamics of bottom-line measures (for example, utility) more a direct reflection of performance rather than an emergent behavior. With regards to setup hypotheses, while our experimental setup parameters are designed for real-world plausibility, the model’s external validity ultimately requires empirical testing. With regards to introspection, we underline that Permutation Importance and SHAP are not meant to reveal a causal relationship between features and outcome, but importance with respect to the predictions computed by the model under consideration. Moreover, we recognize that our definition of utility is focused on economics, thus representing a potentially incomplete picture when seen from another perspective (for example, from the point of view of social sciences). A final consideration pertains the choice of the number of replications $K$ for each simulation, which trades off the processing effort required by the experiment for the estimated margin of error (MOE) related to $\mu(\Lambda^{s})$, $\mu(\Lambda^{s}_{c})$, and $\mu(\Lambda^{s}_{c,d})$ computed over the replication space. While a number of replications between 5000 and 10000 should be recommended (Mundform et al., 2011), one of our goals is to minimize the computational effort to streamline the time and cost of simulations given an acceptable margin of error. Therefore, the considerations based on the analysis of $\mu(\Lambda^{s}_{[c],[d]})$ are conditioned upon the acceptance of the levels of MOE estimated in Section 6, Figures 13 and 13, and in Table 13. For the purpose of our study, we acknowledge the current margin of error as sufficient. Should a lower margin of error be required by the reader, the number of replications $K$ would need to be increased. To this end, our code is available with MIT license upon request.
@@ 8 Conclusions
The common wisdom in the deployment of artificial skills is built on a comforting assumption: combining humans and machines is a defensive safeguard against worst-case scenarios. This policy is believed to mitigate the risks of full automation while capturing some of its benefits by means of a moderate and cautious compromise. This paper formally challenges this assumption, arguing that it may be not only incorrect, but also economically dangerous, thus paradoxical. Our findings reveal that, in high-difficulty tasks, the human-machine skill policy is a high-risk, high-reward strategy. This is the primary contribution of our work. We show that in tasks characterized by high generalization difficulty, when genuine augmentation is not achieved, the human-machine approach frequently becomes the worst of all possible options, and an economic pitfall that actively destroys value by incurring the costs of both human and machine components without a synergistic benefit. In many common scenarios, a manager would be economically better off choosing either a human-only or a machine-only policy. However, the risk is mirrored by an equally high potential reward. When surprise is the norm, if genuine augmentation is achieved, the combination of human and artificial skills can unlock the highest economic utility, outperforming either humans or machines alone. This finding fundamentally highlights the two-folded nature of the managerial challenge of the AI era. The first strategic decision pertains whether or not to blend human and artificial skills. On this subject, our study suggests that the solution to this problem is highly contextual, depending on the skills available, on their performance, and on the difficulty of the task. The second organizational challenge is how to combine human and machine skills towards augmentation. Herein, the default human-in-the-loop paradigm may be insufficient, requiring a focused design effort to develop a collaborative partnership.
The in-silico framework we propose is a direct-to-purpose tool for ex-ante validation, allowing leaders to test the economic viability of their skill policies before costly implementation. By quantifying the economic utility of different skill policies, our experiments provide the foundation for a structured managerial approach. This framework moves beyond static task allocation and instead proposes a dynamic process for effective skill policy evaluation articulated in four stages:
@@
    Economic problem decomposition (understand the puzzle). Before a policy can be evaluated, the managerial problem must be formally decomposed. This requires to i) rigorously define tasks, and classifying them by generalization difficulty, ii) model task-execution performance in terms of distribution of probability for all available skills relevant to the task, and iii) map the economic utility of outcomes, including a clear quantification of the cost of errors;
@@
    Focus on synergistic augmentation (design 1+1=3). Our findings suggest that the human-machine policy is a high-variance strategy whose success is contingent on achieving genuine augmentation. This implies that managers must shift their focus from deciding allocations to designing collaborative workflows that actively engage human and machine skills in a process that can allow a better handling of generalization difficulty;
@@
Ultimately, our framework provides a quantitative method for organizations to move from general heuristics to data-driven policy decisions, allowing for a nuanced test-before-invest evaluation of the trade-offs between automation, human expertise, and human-machine collaboration. Naturally, this study has limitations: while our framework is a simulation grounded in empirical pragmatism, it is not a direct field study. Therefore, empirical testing is the great next step. Additionally, future work could extend this model to include dynamic learning and the impact of different organizational structures. This paper serves as a formal economic challenge to the myth of human-machine integration as a safe harbor. It is a call to move beyond wishful thinking and toward a critical, evidence-based approach for deciding when (and when not) to combine human and artificial skills, improving the organization’s competitiveness and the economic effectiveness of AI initiatives through highly-effective skill policy decisions.
@@ 9 Future developments
Empirical validation would radically strengthen this work, and should be the next-step priority. Additionally, modeling the unitary margins of contribution and the cost of error as probability distributions to reflect the variability of cost structures may be considered. Similarly, exploring other setups for the parameters characterizing the probability distributions of performance may be a promising option, together with considering a variety of classes of distributions. Moreover, the evolution of skills through linear interpolation may be refined to consider the nuances of skill learning and adaptation. Also, modeling the evolution of skills as a non-linear function of cumulative experience, or data generated, is an interesting perspective, in line with the provisions illustrated in (Farboodi and Veldkamp, 2021), where data is a defined as a by-product of production potentially enabling and feeding a feedback loop. Finally, future designs may specifically model dynamic training for machine skills, to account for the differences distinguishing the training of statistical models from heuristics or, more in general, of the models operated by the human mind.

### v6

@@ 5 Discussion
While the structural combined importance of the skill policy $c$ with the augmenting function $a$ in the context of a task presenting generalization difficulty $d$ emerges from the direct analysis of simulation data, it is interesting to better understand the relative importance of the other parameters that take part in the creation of economic value. To this end, we leverage introspection analysis on meta-models learned on our synthetic data by means of i) Permutation Importance analysis and ii) SHapley Additive exPlanations method (SHAP). Permutation Importance analysis (breiman2001random) is a computational method measuring the importance of a feature by determining the increase of the prediction error after the values of that feature are randomly permuted (a larger increase in error denotes a more important feature). In the context of our work, we combine Permutation Importance analysis with cross validation, finally computing mean and standard deviation of the results; also, we aggregate categorical values encoded in one-hot variables in a unique feature, improving interpretability. In contrast, the SHAP method (vstrumbelj2014explaining) (datta2016algorithmic) (lundberg2017unified), named by the acronym of SHapley Additive exPlanations, computes the marginal contribution provided by each feature to the output for any given example, providing local interpretability. Then, global model explanations are obtained by aggregating Shapley values across all examples in the data set, suggesting which features are most influential overall, and how they impact the model’s predictions. As the features $c$ and $a$ have a strong correlation by construction, to avoid compromising the response of permutation analysis and SHAP (strobl2008conditional) (lundberg2017unified), we combine these features in a unique new variable called c_ac\_a, which jointly represents skill policy and augmenting effects.
According to these provisions, we learn the model $\hat{f}:(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})\rightarrow\mu(U_{c,d})$. The absence of correlation between variable is verified by computing the adjusted Generalized Variance Inflation Factor (Adj.GVIF) (fox1992generalized), which results consistently in the range $[1.00,1.01]$ (no correlation). With regards to the statistical learning strategy, we implement a Random Forest regressor (RFr) and a Gradient Boosting regressor (GBr), assessing performance through 10-Fold Cross Validation (metrics: $MSE$ and $R^{2}$, see Table 3). The purpose of measuring the predictive performance is only to make sure that the model is effectively compressing data and learning the information encoded in it, an essential pre-requisite for the interpretation of variable importance.
The results illustrated in Figure 12 and in Figure 13 confirm the remarkable importance of the skill/interaction policy in shaping bottom-line utility, followed by task difficulty. An interesting relative importance is visible for the performance threshold triggering the error shocks ($t_{err}$), the cost of errors ($c_{err}$), the baseline margin of contribution associated to the human-machine skill policy ($mc_{HM}$) and the performance-to-output function $g^{\prime}$, indicating that once the fundamental collaborative dynamic is set, the economic success of the policy is nevertheless influenced by its cost and risk architecture. In contrast, and somehow counter-intuitively, the scalar augmenting factor $\gamma_{HM}$ shows to play a minor role towards utility, suggesting that design of the collaborative process itself (the $a$ function, i.e. the kind of interaction between skills) may be more important than a straightforward performance multiplier. In this sense, a collaborative partnership is something different, not simply something more. When narrowing the meta-learning process to high-difficulty tasks only, the results illustrated in Figures 13(c) and in Figure 13(d) visualize the paradox. The skill/interaction policy c_ac\_a still has the largest impact on utility, but its effect is split: higher-valued red dots (representing the human-machine skill policy with the strongest augmenting effects) drive the highest positive impact on utility, while light-blue dots (representing the non-augmenting human-machine interaction) have a negative impact, confirming that the human-machine skill policy is not a risk-mitigating strategy when the task requires the ability to generalize.
@@
Table 3: Results of 10-Fold Cross Validation over the full set of synthetic data (Exp. 1 + Exp. 2 + Exp. 3)
RFr GBr
Meta-model Avg MSE Avg $R^{2}$ Avg MSE Avg $R^{2}$
$\mu(U_{c,d})=\hat{f}(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$ 0.0071 0.8719 0.0079 0.8587
(a) RFr
(b) GBr
Figure 12: Aggregated Permutation Importance analysis (plots illustrate the mean decrease of $R^{2}$)
(a) RFr
(b) GBr
(c) RFr, $d=High$
(d) GBr, $d=High$
@@
Figure 13: SHAP summary (beeswarm), plots illustrating the mean decrease of $R^{2}$ for the model. Variables are ranked by global importance. Each point represents the SHAP value for a specific instance, with the color indicating the feature’s value. Categorical values (extremes): for task difficulty $d$, blue dots correspond to Low and red dots to High; for skill/interaction policy c_ac\_a, blue dots correspond to $H$ and red dots to $HM++$, with intermediate color scale passing through HMHM (light blue), $M$ (grey), and $HM+$ (light red); for the function $g^{\prime}$, the blue dots represent logistic, while red dots represent inverse_logistic.
@@ 6 Limitations
Combining Monte Carlo simulations with machine learning meta-models can provide valuable insights, but limitations must be acknowledged. The primary limitation of this study is that its conclusions are derived from synthetically generated data. The simulation creates an idealized in silico environment whose potential gaps with the real-world should always be considered with the highest regard, as unmodeled interactions could impact the outcomes (law2007simulation). As a consequence, the validity of our findings is contingent upon the fidelity of our model, which is governed by two sets of assumptions: design hypotheses, related to the theoretical mechanics and interactions encoded in the model, and setup hypotheses, concerning the specific characterization of the scenarios of interest through the chosen parameters and probability distributions.
@@
Design hypotheses include, by instance, the specific choice of $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$, and the logics of computation of the economic measures, including the treatments related to error shocks and their cost. Another arbitrary design choice pertains the specific integration of stochastic effects, which in our work are limited to the task-execution performance, leaving the other measures of economic impact to deterministic computation. This can potentially inject a re-scaling effect down through the funnel of economic value, thus making the dynamics of bottom-line measures (for example, utility) more a direct reflection of performance rather than an emergent behavior. With regards to setup hypotheses, while our experimental setup parameters are designed for real-world plausibility, the model’s external validity ultimately requires empirical testing. As a more general principle, results and insights should be intended specific to our explored parameter space. With regards to introspection, we underline that Permutation Importance and SHAP are not meant to reveal a causal relationship between features and outcome, but importance with respect to the predictions computed by the model under consideration. Moreover, we recognize that our definition of utility is focused on economics, thus representing a potentially incomplete picture when seen from another perspective (for example, from the point of view of social sciences).
A final consideration pertains the choice of the number of replications $K$ for each simulation, which trades off the processing effort required by the experiment for the estimated margin of error (MOE) related to $\mu(\Lambda^{s})$, $\mu(\Lambda^{s}_{c})$, and $\mu(\Lambda^{s}_{c,d})$ computed over the replication space. While a number of replications between 5000 and 10000 should be recommended (mundform2011number), one of our goals is to minimize the computational effort to streamline the time and cost of simulations given an acceptable margin of error. Therefore, the considerations based on the analysis of $\mu(\Lambda^{s}_{[c],[d]})$ are conditioned upon the acceptance of the levels of MOE estimated in Section 5, Figure 11 and Table 2. For the purpose of our study, we acknowledge the current margin of error as sufficient. Should a lower margin of error be required by the reader, the number of replications $K$ would need to be increased. To this end, our code is available with MIT license upon request.
@@ 7 Conclusions
The common wisdom in the deployment of artificial skills is built on a comforting assumption: combining humans and machines is a defensive safeguard against worst-case scenarios. This policy is believed to mitigate the risks of full automation while capturing some of its benefits by means of a moderate and cautious compromise. This paper formally challenges this assumption, arguing that it may be not only incorrect but also economically dangerous, and therefore paradoxical.
Our findings reveal that in high-difficulty tasks, the human-machine skill policy becomes a high-risk, high-reward strategy. This is the primary contribution of our work. Jointly, we show that in tasks characterized by high generalization difficulty, when genuine augmentation is not achieved, the human-machine approach frequently becomes the worst of all possible options, and an economic pitfall that actively destroys value by incurring the costs of both human and machine components without a synergistic benefit. However, the risk is mirrored by an equally high potential reward. When surprise is the norm, if genuine augmentation is achieved, the combination of human and artificial skills can unlock the highest economic utility, outperforming either humans or machines alone.
This finding fundamentally highlights the two-folded nature of the managerial challenge of the AI era. The first strategic decision pertains whether or not to blend human and artificial skills. On this subject, our study suggests that the solution to the problem is highly contextual, depending on the skills available, on their performance, on their cost/margin profile and on the difficulty of the task under consideration. The second organizational challenge pertains how to combine human and machine skills towards augmentation when task difficulty is high, because this is the critical success factor conditioning the bottom-line economic results of skill policy decisions. To this end, the default human-in-the-loop paradigm may be insufficient, requiring a focused design effort to develop a collaborative partnership.
In the context of these two challenges, the in-silico framework we propose is a direct-to-purpose tool for ex-ante validation, allowing leaders to test the economic viability of their strategies before costly implementation. By quantifying the economic utility of the skill policies of interest, our experiments provide the foundation for a structured managerial test-before-invest approach supported by the evaluation of the trade-offs between automation, human expertise, and human-machine collaboration. This framework moves beyond static task allocation and instead proposes a dynamic process for effective skill policy evaluation articulated in four stages:
@@
    Economic problem decomposition (understand the puzzle). Before a policy can be evaluated, the managerial problem must be formally decomposed. This requires to i) rigorously define tasks and to classify them by generalization difficulty, ii) model task-execution performance in terms of distribution of probability for all available skills relevant to the task, and iii) map the economic utility of outcomes, including the cost of errors;
@@
    Focus on synergistic augmentation (design 1+1=3). Our findings suggest that, confronting high-difficulty tasks, the human-machine policy is a high-variance strategy whose success is contingent on achieving genuine augmentation. This implies that managers must shift their focus from deciding allocations to designing collaborative workflows that actively engage human and machine skills in a process that can allow a better handling of generalization difficulty;
@@
Naturally, this study has limitations: while our framework is built with plausibility in mind, it is not a direct field study. Therefore, it should be viewed as a computational thought experiment designed to test the logical consequences of our plausible assumptions, requiring future empirical testing for final validation. Additionally, future work could extend this model to include dynamic learning and the impact of different organizational structures.
This paper serves as a formal economic challenge to the myth of human-machine integration as a safe harbor. It is a call to move beyond wishful thinking and toward a critical, evidence-based approach for deciding when (and when not) to combine human and artificial skills, improving the organization’s competitiveness and the economic effectiveness of AI initiatives through highly-effective skill policy decisions. Finally, we hope this work contributes to seeing a thoughtful combination of human and machine skills not merely as an ethical consideration, but as an economically-justified strategy.
@@ 8 Future developments
Empirical validation would radically strengthen this work, and should be the next-step priority. Additionally, modeling the unitary margins of contribution and the cost of error as probability distributions to reflect the variability of cost structures may be considered. Similarly, exploring other setups for the parameters characterizing the probability distributions of performance may be a promising option, together with considering a variety of classes of distributions. Moreover, the evolution of skills through linear interpolation may be refined to consider the nuances of skill learning and adaptation. Also, modeling the evolution of skills as a non-linear function of cumulative experience, or data generated, is an interesting perspective, in line with the provisions illustrated in (farboodi2021model), where data is a defined as a by-product of production potentially enabling and feeding a feedback loop. Finally, future designs may specifically model dynamic training for machine skills.
@@ 9 Acknowledgments
Appendix A Appendix
Table 4: Statistical characterization of the distributions $\omega(\Lambda^{s}_{c,d})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$ by mean, standard deviation, range and median value (Exp. 1, 2 and 3)
Table 5: Statistical characterization of the distributions $\omega(\Lambda^{s}_{c,d})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$ by mean, standard deviation, range and median value (Exp. 1*, 2* and 3*)
Table 6: Statistical characterization of the distributions $\omega(\Lambda^{s}_{c,d})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$ by mean, standard deviation, range and median value (Exp. 1# and 1£)
(a) $\mu(U^{s}_{c,d})$, Exp. 1
(b) $\mu(U^{s}_{c,d})$, Exp. 2
(c) $\mu(U^{s}_{c,d})$, Exp. 3
(d) $\mu(U^{s}_{c,d})$, Exp. 1*
(e) $\mu(U^{s}_{c,d})$, Exp. 2*
(f) $\mu(U^{s}_{c,d})$, Exp. 3*
(g) $\mu(U^{s}_{c,d})$, Exp. 1#
(h) $\mu(U^{s}_{c,d})$, Exp. 1£
Figure 14: Distributions of the $\mu(U^{s}_{c,d})$, with breakdown by skill policy, task difficulty, and augmenting effect $a$ (overlapping violin and box plots)
(a) $\sigma(U^{s}_{c,d})$, Exp. 1
(b) $\sigma(U^{s}_{c,d})$, Exp. 2
(c) $\sigma(U^{s}_{c,d})$, Exp. 3
(d) $\sigma(U^{s}_{c,d})$, Exp. 1*
(e) $\sigma(U^{s}_{c,d})$, Exp. 2*
(f) $\sigma(U^{s}_{c,d})$, Exp. 3*
(g) $\sigma(U^{s}_{c,d})$, Exp. 1#
(h) $\sigma(U^{s}_{c,d})$, Exp. 1£
Figure 15: Distributions of the $\sigma(U^{s}_{c,d})$, with breakdown by skill policy, task difficulty, and augmenting effect $a$ (overlapping violin and box plots)
Algorithm 2 Function $randomchoice:(X,n,P(X))\rightarrow choice$
[…]
2:$n$
[…]
Return: choicechoice $\triangleright$ A vector of length $n$
[…]
3:$n$
[…]
Return: samplesample $\triangleright$ A vector of length $n$
1:param_startparam\_start
2:param_endparam\_end
3:cur_stepcur\_step
4:tot_stepstot\_steps
5:if $total\_steps\leq 1$ then
6:  $cur\_param\leftarrow param\_start$
7:else
9:end if
Return: cur_paramcur\_param